OpenStudy (anonymous):

holly goes to the store to buy peanut butter and jelly for a school picnic.peanut butter costs$4 a jar,and jelly costs $2 a jar.if holly spends a total of $28 and buys a total of 8 jars,how many jars of peanut butter did she buy? a)12 b)8 c)7 d)6 e)2 ??? btw i looove peanut butter????

OpenStudy (anonymous):

4x+2y=28 x+y=8

OpenStudy (anonymous):

me too love peanut butter 2 <------

OpenStudy (angela210793):

4b+2j=28 b+j=8(*-2) ---> 2b-2j=-16 add both eq 6b=12 b=2--->j=8-2=6

OpenStudy (anonymous):

2<-----peanut butter 6<--jelly